# printacmref=true

### v1

@@ 4. Entropic Exploration for Constrained MARL
The typical goal of constrained MARL is to maximize a reward function while satisfying constraints modeling safe behaviors. Recalling the cost and reward-based objective notations of Sections 2.1 and 3, constrained problems can be defined as maximizing the joint reward objective—$\max_{\pi}J_{r}(\pi)$—while satisfying constraints—$J_{j}^{i}(\pi)\leq l_{j}^{i}\leavevmode\nobreak\ \forall i\in\mathcal{N},\leavevmode\nobreak\ j=1,\dots,m^{i}$ for individual constraints, or $J_{j}(\pi)\leq l_{j}\leavevmode\nobreak\ \forall j=1,\dots,m$ for team ones.
@@
A typical solution to Equation 7 is to iteratively take gradient ascent steps in $\pi$ and descent in $\bm{\lambda}$. We first update the multipliers following $\nabla_{\bm{\lambda}}\mathcal{L}^{\pi}(\bm{\lambda})$, noting the multipliers must be $\geq 0$ because they act as a penalty when the constraints are not satisfied (i.e., $\bm{\lambda}$ increases) while decreasing to 00 (i.e., removing any penalty) when the constraint objectives are satisfied. Then, we maximize the policy’s parameters following $\nabla_{\pi}\mathcal{L}^{\pi}\big{(}\bm{\lambda})$. Broadly speaking, Lagrangian-based algorithms focus on satisfying the imposed constraints using a penalty growing unbounded when constraints are violated. Once the constraints are satisfied, the multipliers scale down (to zero) and allow the gradient to follow the direction that maximizes the main reward objective. Such formalization can be used to optimize arbitrary policy gradient objectives.
@@ 4.1.1. Entropic Exploration
1:Define flags $count\_based$, $knn\_approximation$, and $mix\_with\_extrinsic\_reward$
@@
3:if $count\_based$ then
@@
6:else if $knn\_approximation$ then
@@
12:     $\leftarrow{\beta}(o_{t+1}^{i})\textit{reward}$ reward
@@
14:if $mix\_with\_extrinsic\_reward$ then
15:     $\leftarrow r_{ex}(o_{t+1}^{i})+\psi\textit{reward}$ reward

### v2

@@ 4. Entropic Exploration for Constrained MARL
[…]
@@
A typical solution to Equation 7 is to iteratively take gradient ascent steps in $\pi$ and descent in $\bm{\lambda}$. We first update the multipliers following $\nabla_{\bm{\lambda}}\mathcal{L}^{\pi}(\bm{\lambda})$, noting the multipliers must be $\geq 0$ because they act as a penalty when the constraints are not satisfied (i.e., $\bm{\lambda}$ increases) while decreasing to 0 (i.e., removing any penalty) when the constraint objectives are satisfied. Then, we maximize the policy’s parameters following $\nabla_{\pi}\mathcal{L}^{\pi}\big{(}\bm{\lambda})$. Broadly speaking, Lagrangian-based algorithms focus on satisfying the imposed constraints using a penalty growing unbounded when constraints are violated. Once the constraints are satisfied, the multipliers scale down (to zero) and allow the gradient to follow the direction that maximizes the main reward objective. Such formalization can be used to optimize arbitrary policy gradient objectives.
@@ 4.1.1. Entropic Exploration
1:Define flags count_basedcount\_based, knn_approximationknn\_approximation, and mix_with_extrinsic_rewardmix\_with\_extrinsic\_reward
@@
3:if count_basedcount\_based then
@@
6:else if knn_approximationknn\_approximation then
@@
12:  $\textit{reward}\leftarrow{\beta}(o_{t+1}^{i})\textit{reward}$
@@
14:if mix_with_extrinsic_rewardmix\_with\_extrinsic\_reward then
15:  $\textit{reward}\leftarrow r_{ex}(o_{t+1}^{i})+\psi\textit{reward}$